OpenStudy (anonymous):

An electron in the cathode ray tube of a television set enters a region where it accelerates uniformly from a speed of 59900 m/s to a speed of 4.06 × 10 6 m/s in a distance of 3.83 cm. What is its acceleration? Answer in units of m/s

OpenStudy (anonymous):

you have two points. A=(y,x)=(59900, 0) and B=(y,x)=(4060000, 0.00383) plug each point into y=mx +b and solve for m. m is slope. the slope of the velocity between two points is acceleration. A: 59900=m(0) +b ---> b = 59900 B: 4060000=m(0.00383) + 59900 -----> 4000100 = m(0.00383) -----> m= 4000100/0.00383 = 1044412532.637